KyledaGreat:

Evaluate this function at x=−3. Express your answer as an integer or simplified fraction. If the function is undefined at the given value, indicate "Undefined".

KyledaGreat:

1 attachment
KyledaGreat:

@lowkey

lowkey:

How do you think of approaching this problem now with the stuff we went over last post?

KyledaGreat:

go over the second function

KyledaGreat:

\[\frac{ 1 }{ 2} (-3) + 5 = \frac{ 7 }{ 2}\]

lowkey:

Yes! and the reason why we go over the second function is because the domain is greater than or equal to -3, and x=-3 satisfied this function's domain.

KyledaGreat:

ohhh i see

KyledaGreat:

okay Evaluate this function at x=−7. Express your answer as an integer or simplified fraction. If the function is undefined at the given value, indicate "Undefined".